Caffeine is a weak base with pKa of 14. In the small

intestine, with a pH of 6, what is the ratio of

nonionized to ionized caffeine? Is absorption in the small intestine favored or not and why?
Answer #1

Answer :

Caffeine is a weak base with pKa of 14.

In small intestine a pH of 6.

The ratio of nonionized to ionized caffeine :

pH - pKa = 6 - 14 = -8

Since caffeine is abase drug.

Ratio back words : nonionized / ionized

10 3 / 1 = 1000/ 1

In the small intestine caffeine will be mostly nonionized and it is a fat soluble.

The nonionized molecules pass through intestinal membranes and absorption will be taken place.

So the small intestine is favorable to absorption of caffeine, because it is nonionized and easily pass through intestinal membranes.